# Continuous-Aperture Array (CAPA)-Based Wireless Communications: Capacity Characterization

### v1

@@ IV-B2 Capacity Characterization
Figure 2: Geometries of CAPAs.
V Case Studies
As discussed in Remark 3 and Remark 8, the sum-rate capacity expressions derived for both uplink and downlink CAPA communications are applicable to apertures of arbitrary location, shape, and size, among other factors. In the following sections, we specialize ${\mathcal{A}}$ to several specific cases to unveil more system insights.
@@ V-A Planar CAPAs
We first consider the case where the BS array is a planar CAPA placed on the $x$-$z$ plane and centered at the origin, as shown in Fig. 2(a). The edges of $\mathcal{A}$ are parallel to the axes, with physical dimensions $L_{x}$ and $L_{z}$ along the $x$- and $z$-axes, respectively. In this case, we have ${\mathcal{A}}=\{[x,0,z]^{\mathsf{T}}|x,z\in[-\frac{L_{x}}{2},\frac{L_{x}}{2}]\}$. For each user $k\in\{1,2\}$, let $r_{k}$ represent the distance from the center of $\mathcal{A}$ to the center of $\mathcal{A}_{\mathsf{k}}$, $\phi_{k}\in[0,\pi]$ and $\theta_{k}\in[0,\pi]$ denote the associated azimuth and elevation angles, respectively. Consequently, ${\mathbf{s}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Theta_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\cos{\phi_{k}}\sin{\theta_{k}}$, $\Psi_{k}\triangleq\sin{\phi_{k}}\sin{\theta_{k}}$, and $\Theta_{k}\triangleq\cos{\theta_{k}}$. Based on (8) and (15), for $\mathbf{r}=[x,0,z]^{\mathsf{T}}\in\mathcal{A}$, we have
@@
To characterize the capacity, we derive closed-form expressions of the channel gain and correlation factor as follows.
@@
$\begin{split}\rho=&\frac{\pi^{2}A}{4n^{2}\sqrt{\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}{\sqrt{\left(1-\psi_{j}^{2}\right)\left(1-\psi_{j^{\prime}}^{2}\right)}}\\ &\times\mathsf{Q}_{1}^{*}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\triangleq\rho_{\mathsf{p}},\end{split}$ (61)
@@
To shed more light on the characteristics of planar CAPAs, we investigate the asymptotic sum-rate capacity for a limiting case where $\mathcal{A}$ is assumed to be infinitely large, i.e., $L_{x},L_{z}\rightarrow\infty$. In this case, we have $\lim_{L_{x},L_{z}\rightarrow\infty}{\mathsf{g}}_{k}^{\mathsf{p}}=\frac{1}{4\pi}\frac{4\pi}{2}=\frac{1}{2}$. Regarding $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}$, while it is computationally intractable, based on the numerical results presented in [23, 24], we observe that $\lim_{L_{x},L_{z}\rightarrow\infty}{\rho}_{\mathsf{p}}\ll 1$. Taken together, we can obtain the asymptotic sum-rate capacity as follows:
@@
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. The results in (63) suggest that when the BS is equipped with an infinitely large planar CAPA, half of the transmission power of each user can be captured by the receiver, and the inter-user interference vanishes.
@@ V-B Linear CAPAs
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. By comparing (63) with (66), we observe that the asymptotic sum-rate capacity achieved by the linear CAPA is influenced not only by the transmit SNR but also by the locations of the users, which is in contrast to the planar CAPA.
Figure 3: Planar SPDA.
@@ V-C Planar SPDAs
Finally, we consider a case where the above planar CAPA is partitioned into $M=M_{z}M_{x}$ spatially discrete elements, where $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$ denote the number of antenna elements along the $x$- and $z$-axes, as depicted in Fig. 3. The physical dimensions of each element are indicated by $\sqrt{{A}_{\mathsf{s}}}$, and the inter-element distance is denoted as $d$, where $\sqrt{{A}_{\mathsf{s}}}\leq d\ll r_{k}$. In this case, the central position of each element is given by $[m_{x}d,0,m_{z}d]^{\mathsf{T}}$ for $m_{x}\in\mathcal{M}_{x}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{x}\}$ and $m_{z}\in\mathcal{M}_{z}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{z}\}$. Further, we have $L_{x}\approx M_{x}d$, $L_{z}\approx M_{z}d$, and $\mathcal{A}=\{[m_{x}d+\ell_{x},0,m_{z}d+\ell_{z}]^{\mathsf{T}}|\ell_{x},\ell_{z}\in[-\frac{\sqrt{{A}_{\mathsf{s}}}}{2},\frac{\sqrt{{A}_{\mathsf{s}}}}{2}],m_{x}\in\mathcal{M}_{x},m_{z}\in\mathcal{M}_{z}\}$.
@@ VI Numerical Results
In this section, computer simulations are performed to demonstrate the performance of CAPA communications and verify the accuracy of the developed analytical results. For clarity, the simulations employ the planar arrays specified in Section V. Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.0107$ m, $d=\frac{\lambda}{2}$, $A_{\mathsf{u},1}=A_{\mathsf{u},2}=\frac{\lambda^{2}}{4\pi}$, $L_{x}=L_{z}$, $M_{x}=M_{z}$, $\overline{\gamma}_{\mathsf{ul},1}=30$ dB, $\overline{\gamma}_{\mathsf{ul},2}=40$ dB, $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}=50$ dB with $\sigma_{1}^{2}=\sigma_{2}^{2}$, $\theta_{1}=\theta_{2}=\frac{\pi}{6}$, $\phi_{1}=\phi_{2}=\frac{\pi}{3}$, $r_{1}=10$ m, and $r_{2}=20$ m. All the results are calculated by setting the complexity-vs-accuracy tradeoff parameter $n$ as $n=200$.
Figure 4: Uplink transmission rates achieved by CAPA.
(a) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
(b) Sum-rate capacity vs. $A$.
@@ VI-A Uplink CAPA Communications
Figure 4 illustrates the uplink per-user rate and sum-rate capacity achieved by the CAPA in terms of the BS aperture size under the SIC decoding orders $1\rightarrow 2$ and $2\rightarrow 1$. It can be observed from this graph that as the aperture size $A$ grows, all the presented rates increase. This increase is attributed to the larger aperture size enhancing channel gain and reducing channel correlation, which in turn enhances the dedicated signal and diminishes the IUI, as discussed in (63).
To compare the performance gap between CAPA and the conventional SPDA in terms of sum-rate capacity, we plot the sum-rate capacity for SPDAs as a function of the array occupation ratio $\zeta_{\mathsf{oc}}$ in Figure 5(a). As shown in the graph, the sum-rate capacity for an SPDA gradually converges with that of a CAPA as the array occupation ratio increases towards $1$. This observation validates the statements in Remark 10. Additionally, the results in Figure 5(a) indicate that increasing the aperture size can boost the channel capacity. To gain further insights into the capacity limit as the BS aperture size $A$ approaches infinity, we plot the sum-rate capacity against the aperture size in Figure 5(b). As anticipated, increasing the aperture size causes the sum-rate capacity to approach its upper bound, as derived in (63) and (67). It is worth mentioning that the asymptotic sum-rate capacity for an infinitely large aperture is a finite value, subject to the principle of energy conservation [23]. This verifies our discussions in Remark 9.
@@
(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$.
@@
Figure 6 shows the uplink capacity region achieved by the CAPA, where the green and blue points represent the achieved rates by SIC decoding order $1\rightarrow 2$ and $2\rightarrow 1$, respectively. The rate tuple on the line segment connecting these two points is achieved by the time-sharing strategy. In Figure 6(a), the capacity regions achieved by the CAPA for various aperture sizes are illustrated. It can be observed that the capacity region expands with increasing $A$, shifting from a pentagonal to a rectangular shape. This phenomenon occurs because the IUI decreases as the aperture size increases, which is consistent with the results shown in Figure 4. Furthermore, as depicted in Figure 6(b), the capacity region of the CAPA surpasses that of the SPDA, and their disparity narrows as the array occupation rate $\zeta_{\mathsf{oc}}$ increases, reinforcing findings from Figure 5(a). These findings underscore the superiority of CAPAs over conventional SPDAs in terms of channel capacity.
Figure 7: Downlink transmission rates achieved by CAPA.
@@ VI-B Downlink CAPA Communications
Figure 7 plots the per-user rate and the sum-rate capacity in terms of the total transmit SNR $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}$ in downlink CAPA communications. It is observed that the transmission rates of both users exhibit identical high-SNR slopes, regardless of the DPC encoding orders. Figure 8(a) and Figure 8(b) illustrate the downlink sum-rate capacity with respect to the aperture size and array occupation ratio, respectively. Similar to the uplink case, the downlink sum-rate capacity converges to its upper bound as the aperture size increases. Additionally, for the SPDA, the achieved sum-rate capacity is positively correlated with the array occupation ratio, further suggesting that the CAPA outperforms the SPDA in terms of channel capacity.
(a) CAPA capacity region vs. $A$.
(b) Capacity region vs. $\zeta_{\mathsf{oc}}$, $A\!=\!2\,m^{2}$
Figure 9: Downlink capacity regions.
Figure 9 illustrates the downlink capacity region achieved by the CAPA. The downlink capacity region is depicted by the convex hull of all the corresponding dual uplink capacity regions with the same sum power budget [20]. This means that the boundary of the downlink capacity region is formed by the corner points of the dual uplink capacity regions, as illustrated in Figure 9(a). It is observed that the downlink capacity region of the CAPA expands with the aperture size $A$, whereas the capacity region of the SPDA expands with the array occupation ratio $\zeta_{\mathsf{oc}}$. Taken together, these observations demonstrate that the CAPA achieves a more extensive downlink capacity region than the conventional SPDA, highlighting its superiority.

### v2

@@ V Case Studies
Figure 2: Illustration of the geometries of CAPAs.
@@ V-A Planar CAPAs
We first consider the case where the BS array is a planar CAPA placed on the $x$-$z$ plane and centered at the origin, as shown in Fig. 2(a). The edges of $\mathcal{A}$ are parallel to the axes, with physical dimensions $L_{x}$ and $L_{z}$ along the $x$- and $z$-axes, respectively. In this case, we have $\mathcal{A}=\{[x,0,z]^{\mathsf{T}}|x\in[-\frac{L_{x}}{2},\frac{L_{x}}{2}],z\in[-\frac{L_{z}}{2},\frac{L_{z}}{2}]\}$. For each user $k\in\{1,2\}$, let $r_{k}$ represent the distance from the center of $\mathcal{A}$ to the center of $\mathcal{A}_{\mathsf{k}}$, $\phi_{k}\in[0,\pi]$ and $\theta_{k}\in[0,\pi]$ denote the associated azimuth and elevation angles, respectively. Consequently, ${\mathbf{s}}_{k}=[r_{k}\Phi_{k},r_{k}\Psi_{k},r_{k}\Theta_{k}]^{\mathsf{T}}$, where $\Phi_{k}\triangleq\cos{\phi_{k}}\sin{\theta_{k}}$, $\Psi_{k}\triangleq\sin{\phi_{k}}\sin{\theta_{k}}$, and $\Theta_{k}\triangleq\cos{\theta_{k}}$. Based on (8) and (14), for $\mathbf{r}=[x,0,z]^{\mathsf{T}}\in\mathcal{A}$, we have
@@
To calculate the capacity, we derive closed-form expressions of the channel gain and correlation factor as follows.
@@
$\begin{split}\rho=&\frac{\pi^{2}A}{4n^{2}\sqrt{\mathsf{g}_{1}^{\mathsf{p}}\mathsf{g}_{2}^{\mathsf{p}}}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}{\sqrt{\left(1-\psi_{j}^{2}\right)\big{(}1-\psi_{j^{\prime}}^{2}\big{)}}}\\ &\times\mathsf{Q}_{1}^{*}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\mathsf{Q}_{2}\Big{(}\frac{L_{x}\psi_{j}}{2},\frac{L_{z}\psi_{j^{\prime}}}{2}\Big{)}\triangleq\rho_{\mathsf{p}},\end{split}$ (65)
@@
Note that the expression in (65) is derived using the Chebyshev–Gauss quadrature rule, as per Appendix -F.
@@
To shed more light on the characteristics of planar CAPAs, we investigate the asymptotic sum-rate capacity for a limiting case where $\mathcal{A}$ is assumed to be infinitely large, i.e., $L_{x},L_{z}\rightarrow\infty$. In this case, we have $\lim_{L_{x},L_{z}\rightarrow\infty}{\mathsf{g}}_{k}^{\mathsf{p}}=\frac{1}{4\pi}\frac{4\pi}{2}=\frac{1}{2}$. Regarding $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert$, while it is computationally intractable, based on the numerical results presented in [54, 55], we observe that $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert\ll 1$. Taken together, we can obtain the asymptotic sum-rate capacity as follows:
@@
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. The results in (67) suggest that when the BS is equipped with an infinitely large planar CAPA, half of the transmission power of each user can be captured by the receiver, and the IUI vanishes.
@@
Additionally, since $\lim_{L_{x},L_{z}\rightarrow\infty}\lvert{\rho}_{\mathsf{p}}\rvert\ll 1$, we have
$\lim_{L_{x},L_{z}\rightarrow\infty}\mathsf{R}_{i}^{\mathsf{zf}}\approx\lim_{L_{x},L_{z}\rightarrow\infty}\mathsf{C}_{i},$ (68)
for $i\in\{\mathsf{ul},\mathsf{dl}\}$. This suggest that when the size of the CAPA is sufficiently large, the sum-rates achieved by the ZF schemes can approach the sum-rate capacities.
@@ V-B Linear CAPAs
where $i\in\{\mathsf{ul},\mathsf{dl}\}$. By comparing (67) with (71), it is found the asymptotic sum-rate capacity achieved by the linear CAPA is influenced not only by the SNR but also by the locations of the users, which is in contrast to the planar CAPA.
Figure 3: Illustration of a planar SPDA.
@@ V-C Planar SPDAs
Finally, we consider a planar SPDA with the same size as the planar CAPA, i.e., $L_{x}\times L_{z}$, comprising $M=M_{z}M_{x}$ spatially discrete elements, where $M_{x}=2\tilde{M}_{x}+1$ and $M_{z}=2\tilde{M}_{z}+1$ denote the number of antenna elements along the $x$- and $z$-axes, as depicted in Fig. 3. The physical dimensions of each element are indicated by $\sqrt{{A}_{\mathsf{s}}}$, and the inter-element distance is denoted as $d$, where $\sqrt{{A}_{\mathsf{s}}}\leq d\ll r_{k}$. In this case, the central position of each element is given by $[m_{x}d,0,m_{z}d]^{\mathsf{T}}$ for $m_{x}\in\mathcal{M}_{x}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{x}\}$ and $m_{z}\in\mathcal{M}_{z}\triangleq\{0,\pm 1,\ldots,\pm\tilde{M}_{z}\}$. Further, we have $L_{x}\approx M_{x}d$, $L_{z}\approx M_{z}d$, and $\mathcal{A}=\{[m_{x}d+\ell_{x},0,m_{z}d+\ell_{z}]^{\mathsf{T}}|\ell_{x},\ell_{z}\in[-\frac{\sqrt{{A}_{\mathsf{s}}}}{2},\frac{\sqrt{{A}_{\mathsf{s}}}}{2}],m_{x}\in\mathcal{M}_{x},m_{z}\in\mathcal{M}_{z}\}$.
@@
We also note that the above conclusion is derived without considering the impact of EM coupling. When MC is taken into account, the relationship between the sum-rates achieved by CAPAs and SPDAs may depend on the size of each discrete element as well as the operating frequency [50].
(a) $\left|\rho\right|^{2}$ vs. $n$ for small aperture sizes.
(b) $\left|\rho\right|^{2}$ vs. $n$ with $A=10^{6}$ m2.
Figure 4: Convergence of Chebyshev-Gauss quadrature.
Figure 5: Uplink transmission rates of CAPA.
@@ VI Numerical Results
In this section, computer simulations are performed to demonstrate the performance of CAPA-based communications and verify the accuracy of the developed analytical results. For clarity, the simulations employ the planar arrays specified in Section V. Unless otherwise specified, the simulation parameters are set as follows: $\lambda=0.125$ m, $A_{\mathsf{u},1}=A_{\mathsf{u},2}=A_{\mathsf{s}}=\frac{\lambda^{2}}{4\pi}$, $L_{x}=L_{z}=0.5$ m2, $M_{x}=M_{z}$, $\overline{\gamma}_{\mathsf{ul},1}=30$ dB, $\overline{\gamma}_{\mathsf{ul},2}=40$ dB, $\overline{\gamma}_{\mathsf{dl},1}+\overline{\gamma}_{\mathsf{dl},2}=50$ dB with $\sigma_{1}^{2}=\sigma_{2}^{2}$, $\theta_{1}=\theta_{2}=\frac{\pi}{6}$, $\phi_{1}=\phi_{2}=\frac{\pi}{3}$, $r_{1}=10$ m, and $r_{2}=20$ m.
To select an appropriate value for the complexity-vs-accuracy tradeoff parameter $n$ in the Chebyshev-Gauss quadrature given in (65), we first analyze the convergence behavior of $\left|\rho\right|^{2}$ with respect to $n$ for different aperture sizes, as illustrated in Figure 4. The results indicate that for all aperture sizes depicted in Figure 4(a), $\left|\rho\right|^{2}$ stabilizes once $n$ exceeds $20$. This means that using $20$ terms for summation in (65) can precisely approximate the integral $\int_{{\mathcal{A}}}{\mathsf{G}}_{1}^{*}({\mathbf{r}}){\mathsf{G}}_{2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}$ contained in the channel correlation factor $\left|\rho\right|^{2}=\frac{\left|\int_{{\mathcal{A}}}{\mathsf{G}}_{1}^{*}({\mathbf{r}}){\mathsf{G}}_{2}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\right|^{2}}{{{\mathsf{g}}_{1}{\mathsf{g}}_{2}}}$. Additionally, we observe that larger aperture sizes require higher values of $n$ for the Chebyshev-Gauss quadrature to converge. To ensure convergence for extremely large array sizes, we conduct further simulations by examining an aperture size of $10^{6}$ m2 in Figure 4(b). The results reveal that convergence is effectively achieved when $n=1000$. Based on these observations and to maintain a balance between computational complexity and accuracy, we set $n=1000$ for Figure 6(a) and 9(a) to illustrate asymptotic rate behavior for extremely larger arrays, while using $n=20$ for the other results.
(a) Sum-rate capacity vs. $A$.
(b) Sum-rate capacity vs. $\zeta_{\mathsf{oc}}$.
@@ VI-A Uplink CAPA Communications
Figure 5 illustrates the uplink per-user rate, sum-rate capacity, and the sum-rate achieved by the ZF detector in terms of the BS aperture size. It can be observed from this graph that as the aperture size $A$ grows, all the presented rates increase. This increase is attributed to the larger aperture size enhancing channel gain and reducing channel correlation, which in turn enhances the dedicated signal and diminishes the IUI, as discussed in (67). Furthermore, it is important to note that the sum-rate achieved by the linear ZF detector is lower than the sum-rate capacity achieved by the proposed optimal SIC-based detector.
To gain further insights into the capacity limit as the BS aperture size $A$ approaches infinity, we plot the sum-rate capacity against the array size for both the CAPA and SPDA in Figure 6(a). It is evident that the CAPA outperforms the SPDA for arbitrary array size. As anticipated, increasing the aperture size causes the sum-rate capacity to approach its upper bound, as derived in (67) and (72). It is worth mentioning that the asymptotic sum-rate capacity for an infinitely large aperture is a finite value, subject to the principle of energy conservation [54]. This verifies our discussions in Remark 9. Moreover, as the aperture size of the CAPA grows, the sum-rate achieved by the ZF detector asymptotically approaches the sum-rate capacity, which is consistent with the results in (68). This implies that the linear ZF detector, despite its lower implementation complexity, can deliver performance comparable to the optimal SIC scheme in the large-scale CAPA regime. To compare the performance gap between CAPA and the conventional SPDA in terms of sum-rate capacity, we plot the sum-rate capacity for SPDAs as a function of the array occupation ratio $\zeta_{\mathsf{oc}}$ in Figure 6(b). As shown in the graph, the sum-rate capacity for an SPDA gradually converges with that of a CAPA as the array occupation ratio increases towards $1$. This observation validates the statements in Remark 10.
@@
(b) Capacity regions with MC, $A=1$m2 and $d=\frac{\lambda}{3}$.
@@
Figure 7 shows the uplink capacity region achieved by the CAPA, where the two corner points represent the achieved rates by SIC decoding order $1\rightarrow 2$ and $2\rightarrow 1$. The rate tuple on the line segment connecting these two points is achieved by the time-sharing strategy. In Figure 7(a), the capacity regions achieved by the CAPA for various aperture sizes are illustrated. It can be observed that the capacity region expands with increasing $A$, shifting from a pentagonal to a rectangular shape. This phenomenon occurs because the IUI decreases as the aperture size increases, which is consistent with the results shown in Figure 5.
Figure 8: Downlink transmission rates of CAPA.
@@
To further examine the impact of MC, Figure 7(b) illustrates the rate regions with and without MC for both the CAPA and the SPDA. Since an exact MC model for continuous apertures remains an open research question, we approximate the CAPA as an SPDA with densely packed, edge-to-edge antenna elements [48], where the element size is set as $\sqrt{A_{\mathsf{s}}}=d=\frac{\lambda}{3}$. In this case, the coupling matrix is modeled as $\left(z_{a}+z_{t}\right)\left(\mathbf{Z}+z_{t}\mathbf{I}\right)^{-1}$, where $z_{a}$ represents the antenna impedance, $z_{t}$ denotes the termination impedance of each element, and $\mathbf{Z}$ is the mutual impedance matrix [39]. We set $z_{a}=z_{t}=50~{}\Omega$ and model the mutual impedance elements as $\left[\mathbf{Z}\right]_{i,j}=0.1\frac{\mathrm{e}^{-\mathrm{j}k_{0}d_{ij}}}{d_{ij}^{2}}$, where $d_{ij}$ represents the distance between elements. From Figure 7(b), we observe that MC significantly degrades the performance of both CAPA and SPDA systems. However, despite the performance degradation caused by MC, the capacity region of CAPA remains larger than that of SPDA, highlighting its inherent advantage in terms of channel capacity.
(a) CAPA capacity region vs. $A$.
(b) Capacity regions with MC, $A=1$m2 and $d=\frac{\lambda}{3}$.
Figure 10: Downlink capacity regions.
@@ VI-B Downlink CAPA Communications
Figure 8 plots the per-user rates, the sum-rate capacity, and the sum-rate achieved by the ZF precoder as functions of the aperture size in downlink CAPA communications. A notable performance gap is observed between the sum-rate achieved by the linear ZF precoder and the sum-rate capacity. Figure 9(a) and Figure 9(b) illustrate the downlink sum-rate capacity with respect to the aperture size and array occupation ratio, respectively. Similar to the uplink case, as the aperture size increases, the downlink sum-rate capacity converges to its upper bound, and the sum-rate achieved by the ZF scheme converges to the capacity. Additionally, for the SPDA, the achieved sum-rate capacity is positively correlated with the array occupation ratio, further suggesting that the CAPA outperforms the SPDA in terms of channel capacity..
Figure 10 illustrates the downlink capacity region achieved by the CAPA. The downlink capacity region is depicted by the convex hull of all the corresponding dual uplink capacity regions with the same sum power budget [51]. This means that the boundary of the downlink capacity region is formed by the corner points of the dual uplink capacity regions, as illustrated in Figure 10(a). It is also observed that the downlink capacity region of the CAPA expands with the aperture size $A$. In Fig. 10(b), the MC effects are demonstrated using the same method as in the uplink case. Additionally, in downlink scenario, the CAPA also achieves a more extensive capacity region than the conventional SPDA, highlighting its superiority.